# Bayesian Optimization with Formal Safety Guarantees via Online Conformal Prediction

### v1

@@ VII-C Probabilistic Safety Constraint
We now turn to considering scenarios with observation noise $\sigma^{2}_{q}>0$, and aim at evaluating the performance in terms of probabilistic safety requirement (8) and optimality ratio (3). We set the target reliability level $0.91-\delta=$ 0.9 with target violation rate $\alpha=0.1$ for P-SAFE-BOCP, and with $\alpha=0$ for SAFEOPT in accordance with SAFEOPT’s design. For the latter scheme, we set the “safe” value $B=10||q||_{\kappa^{*}}$, while we consider both the well-specified kernel bandwidth $h=h^{*}=1/1.62$, and the misspecified one $h=1/14.58<h^{*}$, as considered also in the previous set of experiments. For all schemes, the excess violation rate probability in (8) is obtained by averaging over 10,000 realizations.
@@ IX Conclusions
Although in this work we have built on SAFEOPT for the acquisition process, the proposed framework could be generalized directly to any other Safe-BO schemes, such as GOOSE[9]. Other possible extensions include accounting for multiple constraints, as well as taking into account contextual information during the optimization process [29].

### v3

@@ VII-C Probabilistic Safety Constraint
We now turn to considering scenarios with Gaussian observation noise $\sigma^{2}_{q}>0$, and aim at evaluating the performance in terms of probabilistic safety requirement (8) and optimality ratio (3). We set the total number of iterations $T=25$, the target reliability level $0.91-\delta=$ 0.9 with target violation rate $\alpha=0.1$ for P-SAFE-BOCP, and with $\alpha=0$ for SAFEOPT in accordance with SAFEOPT’s design. For the latter scheme, we set the “safe” value $B=10||q||_{\kappa^{*}}$, while we consider both the well-specified kernel bandwidth $h=h^{*}=1/1.62$, and the misspecified one $h=1/14.58<h^{*}$, as considered also in the previous set of experiments. For all schemes, the excess violation rate probability in (8) is obtained by averaging over 10,000 realizations.
@@ IX Conclusions
Although in this work we have built on SAFEOPT for the acquisition process, the proposed framework could be generalized directly to any other Safe-BO schemes, such as GOOSE[9]. Other possible extensions include accounting for multiple constraints, multi-fidelity approximations on objective and constraints [36], as well as taking into account contextual information during the optimization process [37].